# References Matter: Investigating the Impact of Reference Set Variation on Summarization Evaluation

### v2

@@ 4.4 Correlation with Human Judgments
In addition to stability, automatic metrics should correlate with humans. We compute correlations for SummEval and GUMSum, for which we have human judgments, [^4] For SummEval, we use the 161616 models studied by Fabbri et al. (2021); for GUMSum, the four LLMs. at the instance and system level as the number of references $k$ increases. [^5] For example, when considering two references, we consider the sets ($\binom{N}{2}$) of human-written references, where $N$ is the total number of references. We compute the scores using such references as gold standard, and report the mean.
@@
Figure 7 reports the instance-level correlation for SummEval (top) and GUMSum (bottom), respectively, versus number of references. We show ROUGEmax; corresponding figures using ROUGEavg are in Appendix H.
We notice weak-to-no correlation on both datasets. All correlations are generally higher on SummEval (where we consider outputs from the pre-LLM era) than on GUMSum (where we consider LLMs), in accordance to previous work showing that correlation with human judgments decreases as the quality of the outputs improves Peyrard (2019). For SummEval, increasing the number of references consistently leads to better correlation. This effect vanishes on GUMSum, where a larger reference set leads to no effect or slightly lower correlation. For SummEval, BERTScore shows the highest correlation on all dimensions but consistency, for which METEOR and ROUGEavg are a better proxy. Notice how the best metric in terms of correlation with human judgment depends on the considered criterion and the available number of references: BLEURT, for example, typically has low correlation when considering one reference only, performing worse than ROUGE. However, its performance improves when more references are considered, surpassing the scores of n-gram-based metrics.
@@ 5 Conclusions
In this work, we investigate how reference sets impact the reliability of reference-based summarization metrics. Our analysis across three multi-reference datasets reveals that, despite their popularity, token-matching metrics such as ROUGE are highly sensitive to the reference(s). This sensitivity leads to instability in system rankings, particularly when only a small number of references are available, which is typical in summarization datasets. In these situations, we thus recommend avoiding such metrics, echoing earlier calls for caution Schmidtova et al. (2024), in favor of model-based or reference-free alternatives.
We find that increasing the number of reference summaries consistently improves both the stability of metric scores and their alignment with human judgments. In these conditions, n-gram-based metrics such as ROUGE-N become more reliable.
Our findings highlight the need to incorporate reference set variation into evaluation frameworks. Future metric development should explicitly account for this variation. We thus advocate for the creation of larger, more diverse multi-reference datasets, as well as for metric designs that are inherently robust to reference variability. Such efforts will be key to ensuring fairer and more reliable and human-aligned evaluation practices in summarization in the LLM era and beyond.
@@
While we assess metric robustness using four systems (all of which are LLMs as opposed to the case for SummEval, where 161616 pre-LLM supervised systems are available) on the GUMSum and DUC2004 datasets, the relatively small number of models limits the generalization of the conclusions we can draw about system-level stability. Future work should include a larger and more diverse set of systems to better assess generalizability.
@@ Appendix A LLM-generated Summaries
For each source, we generate a single summary using each of the four LLMs. For SummEval, the sources corresponding to the 100100100 multi-reference summaries are taken into account. For DUC2004 Task 1, we generate summaries for the whole dataset (489489489 instances). For GUMSum, we focus on the dev and test set (in total 484848 instances). Thus, we generate a total of 254825482548 summaries. We comply with the license of the existing datasets. For newly collected model outputs and human judgments, we follow the license of the corresponding underlying datasets. The codebase will be made publicly available upon publication.
@@ Appendix B Reference-based Metrics
We use the original Google version of BLEURT. [^12] https://github.com/google-research/bleurt We used the recommended checkpoint,131313available at https://storage.googleapis.com/bleurt-oss-21/BLEURT-20.zip which we did not fine-tune.
@@ Appendix E Quality of Scraped References
Figure 9: Models and CNN/DM reference compared to 101010 crowd-sourced references in SummEval. SummEval models are in lighter blue, LLMs in darker blue. The red line indicates the CNN/DM reference score, and the black dotted line represents a random baseline using a different news article as the hypothesis.
@@ Appendix G Model Ranking
Table 4: Rank stability with a single reference. We ranked systems 100100100 times and compute the Kendall tau correlation among such rankings. R, MTR, BS, and BLRT are short for ROUGE, METEOR, BERTScore, and BLEURT respectively.

### v3

@@ 4.4 Correlation with Human Judgments
[…]
@@
Figure 7 reports the instance-level correlation for SummEval (top) and GUMSum (bottom), respectively, versus the number of references. We show ROUGEmax; corresponding figures using ROUGEavg are in Appendix H.
We notice weak-to-no correlation on both datasets. All correlations are generally higher on SummEval (where we consider outputs from the pre-LLM era) than on GUMSum (where we consider LLMs), in accordance to previous work showing that correlation with human judgments decreases as the quality of the outputs improves Peyrard (2019). Additionally, reference-based evaluation itself could be problematic for very high-quality outputs when references are of worse quality than outputs (see Table 2, where model outputs are often scored on par with, or higher than, references) as also argued by Goyal et al. (2023); metrics might also not be sensitive enough for outputs with more similar quality. The observed low correlation could be motivated by the low IAA in the GUMSum human judgments.
For SummEval, increasing the number of references consistently leads to better correlation. This effect vanishes on GUMSum, where a larger reference set leads to no effect or slightly lower correlation. For SummEval, BERTScore shows the highest correlation on all dimensions but consistency, for which METEOR and ROUGEavg are better proxies. Notice how the best metric in terms of correlation with human judgment depends on the considered criterion and the available number of references: BLEURT, for example, typically has low correlation when considering one reference only, performing worse than ROUGE. However, its performance improves when more references are considered, surpassing the scores of n-gram-based metrics.
@@ 5 Conclusions
In this work, we have investigated how reference sets impact the reliability of reference-based summarization metrics. Our analysis across three multi-reference datasets reveals that, despite their popularity, token-matching metrics such as ROUGE are highly sensitive to the reference(s). This sensitivity leads to instability in system rankings, particularly when only a small number of references are available, which is typical in summarization datasets. In these situations, we thus recommend avoiding such metrics, echoing earlier calls for caution Schmidtova et al. (2024), in favor of model-based or reference-free alternatives.
We find that increasing the number of reference summaries consistently improves both the stability of metric scores and their alignment with human judgments. This might be explained by the possibility of representing a larger human diversity from the one hand, and from that of limiting annotator bias on the other. In these conditions, n-gram-based metrics such as ROUGE-N become more reliable.
Our findings highlight the need to incorporate reference set variation into evaluation frameworks. Future metric development should explicitly account for this variation. While we have not specifically studied the dimensions of diversity in human references besides their lexical, syntactical, and semantic variation from Giulianelli et al. (2023), we believe this is an important area of investigation. We speculate that the main challenge would be to collect a reference set that is “diverse enough” to represent human production for a fixed number of references. Future work in this direction needs to identify and characterize the relevant dimensions of diversity. These dimensions might be lexical, stylistic, intentional, or even sociolinguistic (cutting across the previously mentioned dimensions, Grieve et al. 2025). How to collect such references and whether specific guidelines should be adopted is also an open problem. Future work should also explore the role of genre.
We thus advocate for the creation of larger, more diverse multi-reference datasets, as well as for metric designs that are inherently robust to reference variability. Such efforts will be key to ensuring fairer and more reliable and human-aligned evaluation practices in summarization in the LLM era and beyond.
@@
While we assess metric robustness using four systems (all of which are LLMs as opposed to the case for SummEval, where $16$ pre-LLM supervised systems are available) on the GUMSum and DUC2004 datasets, the relatively small number of models limits the generalization of the conclusions we can draw about system-level stability. Future work should include a larger and more diverse set of systems to better assess generalizability.
@@
Acknowledgments
We would like to thank the members of the MaiNLP lab for their valuable feedback, especially to Andreas Säuberli, Diego Frassinelli, Shijia Zhou, Soh-Eun (Ryan) Shim, and Domenico De Cristofaro. We also thank Pingjun Hong, Yuchen Mao, and Sebastian Loftus for their annotations. We recognize the support for Yang Janet Liu [^7] Work was carried out while at MaiNLP, LMU Munich. and Barbara Plank through the ERC Consolidator Grant DIALECT 101043235.
@@ Appendix A LLM-generated Summaries
For each source, we generate a single summary using each of the four LLMs. For SummEval, the sources corresponding to the $100$ multi-reference summaries are taken into account. For DUC2004 Task 1, we generate summaries for the whole dataset ($489$ instances). For GUMSum, we focus on the dev and test set (in total $48$ instances). Thus, we generate a total of $2548$ summaries. We comply with the license of the existing datasets. For newly collected model outputs and human judgments, we follow the license of the corresponding underlying datasets. The codebase will be made publicly available upon publication.
@@ Appendix B Reference-based Metrics
We use the original Google version of BLEURT. [^14] https://github.com/google-research/bleurt We used the recommended checkpoint,151515available at https://storage.googleapis.com/bleurt-oss-21/BLEURT-20.zip which we did not fine-tune.
@@ Appendix E Quality of Scraped References
Figure 9: Models and CNN/DM reference compared to $10$ crowd-sourced references in SummEval. SummEval models are in lighter blue, LLMs in darker blue. The red line indicates the CNN/DM reference score, and the black dotted line represents a random baseline using a different news article as the hypothesis.
@@ Appendix G Model Ranking
Table 4: Rank stability with a single reference. We ranked systems $100$ times and compute the Kendall tau correlation among such rankings. R, MTR, BS, and BLRT are short for ROUGE, METEOR, BERTScore, and BLEURT respectively.